# AiSAQ: All-in-Storage ANNS with Product Quantization for DRAM-free Information Retrieval

### v1

@@ Abstract.
In approximate nearest neighbor search (ANNS) methods based on approximate proximity graphs, DiskANN achieves good recall-speed balance for large-scale datasets using both of RAM and storage. Despite it claims to save memory usage by loading compressed vectors by product quantization (PQ), its memory usage increases in proportion to the scale of datasets. In this paper, we propose All-in-Storage ANNS with Product Quantization (AiSAQ), which offloads the compressed vectors to storage. Our method achieves $\sim$10 MB memory usage in query search even with billion-scale datasets with minor performance degradation. AiSAQ also reduces the index load time before query search, which enables the index switch between muitiple billion-scale datasets and significantly enhances the flexibility of retrieval-augmented generation (RAG). This method is applicable to all graph-based ANNS algorithms and can be combined with higher-spec ANNS methods in the future.
@@ 1. Introduction
In the recent trend of Large Language Models (LLM), possible implementations of Retrieval-Augmented Generation (RAG)(lewis2020retrieval, ) use vector datasets from different domains as their external knowledge sources to generate more accurate answers, which need ANNS methods to retrieve information from those sources. For further improvement of LLMs with RAG, ANNS methods need to switch between multiple indices of different domains corresponding to users’ or applications’ requests. This requires the index data of all sources to be retained in RAM during the service time or loaded from storage every time of a request.
@@
In this paper, we propose AiSAQ: All-in-Storage ANNS with Product Quantization, a novel method of index data placement. Our implementation achieved only $\sim$10 MB of DRAM usage regardless of the scale of datasets. Our optimization for data placement in the storage index brings minor latency degradation compared to DiskANN, despite larger portion of the index are sitting in the storage. Moreover, index switch between multiple billion-scale datasets can be performed in millisecond-order.
@@
  • Thanks to data placement optimization on SSD, AiSAQ achieves millisecond-order latency for $>$95% 1-recall@1 with tiny memory usage.
@@
As a concurrent work, LM-DiskANN(pan2023lm, ) was published just before this paper’s submission. However, our work proves the effectiveness on billion-scale datasets and proposes a new metric called index switching time, both of which are not stated in their paper.
@@ 2. Preliminaries
2.1. Graph-Based ANNS algorithm
@@ 2.2. DiskANN
Algorithm 1 DiskANN Beam Search with re-ranking
@@ 4.1. Datasets and Experimental Conditions
Table 1. Datasets used for AiSAQ experiments
@@
Distance metric Euclid Euclid cosine similarity
@@
We tested query search experiments on an AWS EC2 i4i.8xlarge instance which has 2x 3750 GB instance stores tied up by RAID 0.
@@ 4.3. Query Search Time
(a) SIFT1M
(b) SIFT1B
(c) KILT E5 22M
Figure 3. Recall vs latency graphs of DiskANN and AiSAQ
@@ 5. Conclusion and Future Works
In this paper, we proposed AiSAQ, a novel method of ANNS data offloading to storage. Our method achieved billion-scale ANNS with only $\sim$10 MB memory usage without significant degradation of performance. AiSAQ will be applicable to all graph-based ANNS algorithms, including future higher-spec algorithms.

### v2

@@ Abstract.
Graph-based approximate nearest neighbor search (ANNS) algorithms work effectively against large-scale vector retrieval. Among such methods, DiskANN achieves good recall-speed tradeoffs using both DRAM and storage. DiskANN adopts product quantization (PQ) to reduce memory usage, which is still proportional to the scale of datasets. In this paper, we propose All-in-Storage ANNS with Product Quantization (AiSAQ ), which offloads compressed vectors to the SSD index. Our method achieves $\sim$10 MB memory usage in query search with billion-scale datasets without critical latency degradation. AiSAQ also reduces the index load time for query search preparation, which enables fast switch between muitiple billion-scale indices.This method can be applied to retrievers of retrieval-augmented generation (RAG) and be scaled out with multiple-server systems for emerging datasets. Our DiskANN-based implementation [^1] https://github.com/KioxiaAmerica/aisaq-diskann is available on GitHub.
@@ 1. Introduction
In the recent trend of Large Language Models (LLM), possible implementations of Retrieval-Augmented Generation (RAG)(lewis2020retrieval, ) use vector datasets as their external knowledge sources to generate more accurate answers. In the original paper, Dense Passage Retrieval(dpr2020, ) is used as a “retriever” for seaching the external sources. That retriever uses HNSW(malkov2018hnsw, ) for searching vector data of external sources, and it can be replaced with other suitable ANNS algorithms depending on the demand of applications. In some situations that a RAG chain requires multiple sources, the retriever need to “switch” between indices. This procedure requires the index data of all sources to be retained in RAM during the service time or loaded from storage every time of a request.
@@
In this paper, we propose AiSAQ : All-in-Storage ANNS with Product Quantization , a novel method of index data placement. Our implementation achieved only $\sim$10 MB of DRAM usage regardless of the scale of datasets. Our optimization for data placement in the storage index brings minor latency degradation compared to DiskANN, despite larger portion of the index are sitting in the storage. Moreover, index switch between multiple billion-scale datasets can be performed in millisecond-order and makes cost advantage in multiple-server environments.
@@
  • Thanks to data placement optimization on SSD, AiSAQ achieves millisecond-order latency for ¿95% 1-recall@1 with tiny memory usage.
@@
  • Multiple-server search for large-scale indices makes cost advantage of total DRAM and SSD consumption.
As a concurrent work, LM-DiskANN(pan2023lm, ) was published before this paper’s submission. However, our work proves the scalability mentioning billion-scale datasets and multiple-server systems for emerging size of datasets in the future. We also propose a new metric called index switching time, whose result reinforces the scalability for larger datasets. Both of these points are not stated in their paper.
@@ 2. Preliminaries
2.1. Graph-Based ANNS Algorithms
@@
2.2. Index Switch for Multiple Sources
Some RAG systems retrieve information from external sources using ANNS methods. For further improvement of LLMs with RAG, ANNS methods need to “switch” between multiple indices of different domains corresponding to users’ or applications’ requests. For example, an LLM which generates latest news across multiple fields such as economy, politics and enternainment needs to generate a single answer with multiple sources of those fields. Another example is an LLM chain which uses multiple sources. Represented by LangChain(Chase_LangChain_2022, ), it is common to use a pipeline with mutiple-time sequential LLM inferences which uses the output from an LLM inference as the input for the next inference. In such situations, it is possible that a RAG inference requires a different source than the one used in the preceding inference. Existing ANNS methods have to retain all of the indices on RAM or load data of an index every time from storage to switch between multiple indices. This leads to large DRAM usage or long index load time.
@@ 2.3. DiskANN
[…]
A node chunk of DiskANN is composed of full-precision vector, number and IDs of outneighbors. Node chunks are aligned in blocks.
@@ 4.1. Datasets and Experimental Conditions
Table 1. Datasets and index build parameters used for AiSAQ experiments
@@
Distance metric Euclid Euclid MIPS
Maximum outdegree $R$ 56 52 69
# of PQ subvectors $b_{PQ}$ 128 32 128
@@
We conducted single-server query search experiments (4.2 to 4.4) on an AWS EC2 i4i.8xlarge instance which has 2x 3750 GB instance stores tied up by RAID 0. The environment of multiple-server experiments is detailed in 4.5.
@@ 4.3. Query Search Time
In the high-recall area, AiSAQ keeps $\sim$11 MB memory usage for all datasets, while DiskANN suffers lower QPS to keep recall@1 when memory usage is reduced.
@@ 4.4. Index Switch
4.5. Cost Analysis and Multiple-Server System
Figure 5. Concept of multiple-server AiSAQ system
Figure 6. Cost estimation of DiskANN and AiSAQ (SIFT1B)
Table 5. DiskANN and AiSAQ comparison with 6 containers search (SIFT1B)
DiskANN AiSAQ
Total memory usage (MB) 183,405 66
Average index load time (ms) 14319.25 1.7
Estimated DRAM & SSD cost (USD) 344 103
AiSAQ achieves similar high performance to DiskANN with extremely small memory usage and index load time, which results in inflated SSD index size. To retain the PQ vector information, an AiSAQ index requires $RNb_{PQ}$ bytes of storage, while DiskANN requires $Nb_{PQ}$ bytes of DRAM. Maximum outdegree of the graph $R$ is typically set around 60, while the cost per bit of DRAM is approximately 30 times of that of SSD (DRAM: 1.8 USD/GB vs SSD: 0.054 USD/GB, obtained from DRAMeXchange.com(dramexchange, )). This means AiSAQ is not more cost effective than DiskANN in a single server.
However, it is possible that query search scales out with multiple servers like Figure 5 to handle increased search requests and larger indices. Such situations makes cost advantages to AiSAQ . Figure 6 shows estimated DRAM and SSD costs of DiskANN and AiSAQ . DiskANN search using $n$ servers requires $n$ times of DRAM capacity, while AiSAQ takes only the same capacity of storage as the case with a single server. AiSAQ overtakes DiskANN in terms of the resource cost in more than 2 search servers, which is reasonable for multiple-server search for larger-scale distributed ANNS system.
Table 5 shows experimental results and cost estimation of DiskANN and AiSAQ on multiple-server system. We used Lustre distributed filesystem with 3 storage servers for the index storage. AiSAQ search programs are executed on 6 Docker containers each regarded as a server on a single physical machine with an EPYC 7402P CPU and 224 GB of DRAM. All machines are connected by RDMA over 200 Gb ethernet. DiskANN with 6 containers consumes much more DRAM than AiSAQ , which can be the main factor of total cost. AiSAQ resulted in slower index load time than experiment with direct-attached SSD because of the slower latency over ethernet, but still reasonablly fast time, while DiskANN did not acheive practical load time.
@@ 5. Conclusion and Future Works
In this paper, we proposed AiSAQ , a novel method of ANNS data offloading to storage. Our method achieved billion-scale ANNS with only $\sim$10 MB memory usage without significant degradation of performance. AiSAQ will be applicable to all graph-based ANNS algorithms, and multiple-server system with lower costs.